Use RESCA:
Lulu is three yrs. older than Pearl. In 4 yrs. time, Lulu will be twice as old as Pearl was 3 yrs. ago. Find pearl's age.

Answer by ankor@dixie-net.com(22740)

Lulu is three yrs. older than Pearl.
In 4 yrs. time, Lulu will be twice as old as Pearl was 3 yrs. ago.
Find pearl's age.
:
Write an equation for each statement:
"Lulu is three yrs. older than Pearl."
L = P + 3
:
" In 4 yrs. time, Lulu will be twice as old as Pearl was 3 yrs. ago."
L + 4 = 2(P - 3)
L + 4 = 2P - 6
L = 2P - 6 - 4
L = 2P - 10
From the 1st statement; replace L with (P+3)
P + 3 = 2P - 10
3 + 10 = 2P - P
13 = P
Pearl is 13 yrs old, obviously Lulu is 16
;
:
Check solution in the statement:
"In 4 yrs. time, Lulu will be twice as old as Pearl was 3 yrs. ago."
16 + 4 = 2(13 - 3)
20 = 2(10)
:
I have no idea what RESCA is!